The line,L9,passes through the point (1,-4)and is parallel to the line y + x - 3 = 0, find the equation of L9
Answer by mananth(16946)

PARALLEL


1 x + 1 y = 3
Find the slope of this line
make y the subject
1 y = -1 x + 3
Divide by 1
y = -1 x + 3
Compare this equation with y=mx+b
slope m = -1
The slope of a line parallel to the above line will be the same
The slope of the required line will be -1
m= -1 ,point ( 1 , -4 )
Find b by plugging the values of m & the point in
y=mx+b
-4 = -1 + b
b= -3
m= -1
Plug value of the slope and b in y = mx +b
The required equation is y = - x -3

m.ananth@hotmail.ca